OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations. 2x-y=3 9x-4y=15 the solutions is ? (type an ordered pair)

OpenStudy (campbell_st):

take the 1st equation and rewrite it y = 2x - 3 substitute it into the 2nd equation 9x - 4(2x -3) = 15 all you need to do is distribute, collect like terms and solve for x

OpenStudy (anonymous):

its confusing to me im not understanding it at all.

OpenStudy (anonymous):

multiply equation 1 by 9, and equation 2 by 2, so as to get : \[ 18x - 9y = 27\] \[18x-8y = 30\] subtract 2 from 1 so as to get \[-y = -3 \] so the answer is y =3, and x = (3 + y) /2 = 3, answer is (3,3)

OpenStudy (campbell_st):

oh well no need to explain...@kjos has given you an answer... good luck

OpenStudy (anonymous):

the idea behind solving such problems is to eliminate unknowns one by one

OpenStudy (campbell_st):

the only problem is he has used elimination

OpenStudy (campbell_st):

you require substitution

OpenStudy (anonymous):

thank you i will have to study up more on this type of stuff

OpenStudy (anonymous):

i agree. campbell's solution is the right one. i missed the requirement for substiution.

OpenStudy (anonymous):

so whats the substitution part

OpenStudy (campbell_st):

@kjos the aim of the site it to help understanding... not give answers

OpenStudy (anonymous):

answer help me see how it is solved

OpenStudy (anonymous):

in a weird way, i have not ;-) . Substitution in this case means that we have been given 2 unknowns, x and y. If we can somehow express x in terms of y, or vice versa, we can replace where-ever we see x in terms of this new expression. thus we "substitute" x with this new expression for x. so the substitution method. In this case, we have been given that \[2*x - y =3\] so \[y = 2*x -3\], and then we can "substitute" this whereever we see y e.g. in the second equation

OpenStudy (anonymous):

so there is is another pair of numbers?

OpenStudy (anonymous):

nope, the answer is right, just that there are at least 2 ways of arriving at the solution. The solution i mentioned in the beginning is called the "elimination method", - what you want is the "substitution method", which is what campbell described and i just described (my 2nd post)..

OpenStudy (anonymous):

I got to as far as 9x+-4(-3)+-4(2x)=15

OpenStudy (anonymous):

yep, carry on .... right direction. get all the x's together and the non-x's together

OpenStudy (anonymous):

ok i think i am somewhat understanding it.

OpenStudy (anonymous):

thank you for helping me both of yall.